# D2ST-Adapter: Disentangled-and-Deformable Spatio-Temporal Adapter for Few-shot Action Recognition

### v3

@@ Abstract
Adapting large pre-trained image models to few-shot action recognition has proven to be an effective and efficient strategy for learning robust feature extractors, which is essential for few-shot learning. Typical fine-tuning based adaptation paradigm is prone to overfitting in the few-shot learning scenarios and offers little modeling flexibility for learning temporal features in video data. In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter), which is a novel adapter tuning framework well-suited for few-shot action recognition due to lightweight design and low parameter-learning overhead. It is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. In particular, we devise the anisotropic Deformable Spatio-Temporal Attention module as the core component of D2ST-Adapter, which can be tailored with anisotropic sampling densities along spatial and temporal domains to learn spatial and temporal features specifically in corresponding pathways, allowing our D2ST-Adapter to encode features in a global view in 3D spatio-temporal space while maintaining a lightweight design. Extensive experiments with instantiations of our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods for few-shot action recognition. Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition.
Keywords:
Action recognition Few-shot learning Adapter tuning
footnotetext: Code is available at https://github.com/qizhongtan/D2ST-Adapter.
Figure 1: Comparison with previous adapter-based action recognition methods. Both AIM [51] and DUALPATH [26] pre-learn the spatial and temporal features separately by duplicating the MSA modules, and insert vanilla adapter into each duplicated module for specific feature adaptation. ST-Adapter [25] first attempts to design the specialized adapter for video data using 3D convolution. Unlike these models, our D2ST-Adapter is designed in a dual-pathway architecture to encode spatial and temporal features in a disentangled manner. Furthermore, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the essential component, which configures different sampling density in the spatial and temporal domains to model two pathways specifically and enables D2ST-Adapter to perform feature adaptation in a global view in 3D spatio-temporal space while maintaining a lightweight design.
@@ 1 Introduction
Few-shot action recognition aims to learn an action recognition model from a set of base classes of video data, which has the capability to recognize novel categories of actions using only a few support samples. To this end, learning an effective feature extractor that is generalizable across different classes is crucial to the task. A typical way [2, 54, 1, 44, 42] is to leverage pre-trained large vision models [13, 37, 29] for feature learning by task adaptation.
@@
Most existing methods [28, 43, 42] seek to adapt large vision models to few-shot action recognition by fine-tuning either the entire or partial model. While such methods have achieved promising performance, there are two important limitations that hamper them from exploiting the full potential of the pre-trained models. First, fully fine-tuning the pre-trained models is prone to overfitting in the few-shot learning scenarios whilst partial fine-tuning acquires limited effectiveness due to limited capacity for adaptation. Second, most large vision models are trained on image data due to the difficulty of collecting a large scale of video data. Thus, these fine-tuning based methods require to learn the temporal features as a post-process after feature learning by the pre-trained backbone, typically by constructing an auxiliary module to model the temporal dynamics [36, 46, 44, 43], which has limited effectiveness compared to the learning manner throughout the whole feature encoding stage.
@@
As a prominent parameter-efficient fine-tuning technique [14, 58, 17, 12], adapter tuning [14, 4, 41] embeds learnable adapters with a negligible amount of parameters into pre-trained models flexibly while keeping the model frozen, conducting model adaptation efficiently and effectively. Compared to the fine-tuning adaptation paradigm, adapter tuning is particularly better-suited to few-shot learning scenarios. A straightforward way to apply adapter tuning to image-to-video transfer learning is to employ the vanilla adapter for image data directly to adapt the pre-learned spatial features and temporal features separately, which is exactly adopted by AIM [51] and DUALPATH [26]. As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules either in cascaded (AIM) or parallel (DUALPATH) manners, and insert vanilla adapter in each duplicated module for specific feature adaptation. However, such methods inevitably incur heavyweight model architecture and massive computation overhead. ST-Adapter [25] first attempts to design the specialized adapter for video data by inserting a 3D convolution layer into the vanilla adapter for learning spatio-temporal features while maintaining a lightweight structure. Nevertheless, we investigate two potential limitations of ST-Adapter. First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in the low-data scenarios, such a joint learning paradigm of spatio-temporal features for video data is inferior to learning the spatial and temporal features in a disentangled way, such as SlowFast [8] or two-stream design [31, 40]. Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
In this work we design the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations. Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in  Figure 1, in which the spatial pathway is responsible for capturing the spatial appearance features while the temporal pathway focuses on learning the temporal dynamics. Thus, our model can encode the spatio-temporal features in a disentangled manner. Moreover, we devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA) module as an essential component to model both the spatial and temporal pathways. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space, which first samples a group of reference points in the spatio-temporal space and then learns the offset for each point to shift them to more informative regions in the 3D space. These shifted points are used as key and value pairs for aDSTA to perform sparse self-attention, which allows our D2ST-Adapter to perform feature adaptation in a global view while maintaining a lightweight design.
@@
A novel design of our aDSTA is that we make the sampling density anisotropic along the spatial and temporal domains by specific configuration, which allows for tailoring specialized versions of aDSTA to model the spatial and temporal pathways separately. In particular, we tailor aDSTA-T with denser sampling along the temporal domain than the spatial domain for the temporal pathway since it focuses on capturing temporal features. In contrast, the tailored aDSTA-S for the spatial pathway samples denser reference points along the spatial domain. To conclude, we make the following contributions:
@@
  • We propose D2ST-Adapter, a novel adapter tuning method for few-shot action recognition, which is designed in a dual-pathway architecture to adaptively encode the spatial and temporal features in a disentangled manner. It can adapt large vision models pre-trained on image data to video data efficiently and effectively in few-shot learning scenarios due to the lightweight design.
@@
  • We devise the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, allowing our D2ST-Adapter to encode features in a global view while maintaining a lightweight design.
@@ 2 Related Work
Few-shot image classification. Existing works for few-shot image classification can be categorized into three groups: augmentation-based, optimization-based, and metric-based. Augmentation-based methods mainly hallucinate [45, 11] and generate [20, 55] new training samples adversarially to prevent the model from over-fitting. Optimization-based algorithms [9, 16, 30] aim to train a model that generalizes well and can be quickly adapted to new tasks through a few gradient update steps. Metric-based approaches recognize novel samples through nearest neighbor classification using a specific distance metric, such as cosine similarity [39, 52], Euclidean distance [32, 6, 53], and learnable metrics [34]. Our work follows the metric-based category due to its simplicity and effectiveness, and we focus on the more challenging video setting.
Few-shot action recognition. Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task. The first one is to investigate the spatio-temporal modeling [59, 60, 54, 27, 46, 56, 36, 43, 50, 47]. For example, MTFAN [46] proposes a motion encoder to learn global motion patterns and injects them into each video representations using a motion modulator. STRM [36] enriches local patch features and global frame features for joint spatio-temporal modeling. MoLo [43] designs a motion autodecoder to explicitly extract motion dynamics in a unified network. Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique. Another direction focuses on designing effective metric learning strategy [2, 1, 28, 44, 46, 15, 24, 57]. OTAM [2] utilizes the DTW [23] algorithm to calculate video distances with strict temporal alignment. TRX [28] exhaustively enumerates all sub-sequences of support and query videos and matches them using an attention mechanism. HyRSM [44] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints. We evaluate our D2ST-Adapter using the above three classical matching strategies to demonstrate its effectiveness and robustness.
Adapter tuning. As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [14], and quickly draws attention in many other research areas [4, 41, 12, 35]. A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model. During training, only the parameters of inserted adapters are updated while the original model remains frozen, leading to efficient task adaptation. Recently, some works apply this method to adapt image models for action recognition. AIM [51] duplicates multi-head self-attention modules and plugs adapters after them to separately learn the spatial and temporal features in a cascaded manner. Similarly, DUALPATH [26] explicitly builds the two-stream architecture upon ViT and utilizes adapters to parallelly learn the spatial and temporal features. ST-Adapter [25] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability. Our D2ST-Adapter follows the basic framework of ST-Adapter, and meanwhile optimize the essential technical designs for few-shot action recognition.
@@ 3.1 Overview
Problem Formulation of Few-shot Action Recognition. The aim of few-shot action recognition is to learn an action recognizer from a set of base classes, which is able to recognize novel classes using only a few labeled samples. The task typically follows episodic paradigm [39, 2, 44, 43], which is consistent with general application scenarios for few-shot learning. An episode includes a support set $\mathcal{S}$ consisting of $N$ classes and $K$ labeled samples for each class (referred to as the $N$-way $K$-shot task), as well as a query set $\mathcal{Q}$ that contains unlabeled samples to be classified. In each episode, we aim to classify every query into one of the $N$ classes with the guidance of the support set. Such episodic task setting is consistently adopted during all the training, validation, and test stages.
Adapter Tuning Framework. We design a plug-and-play and lightweight adapter, dubbed D2ST-Adapter, which can be integrated into most existing large vision models. Thus we can leverage the powerful feature encoding capability of pre-trained large vision models by efficient image-to-video model adaptation while incurring only a small amount of parameter-tuning overhead. It is particularly well-suited to few-shot learning scenarios.
Figure 2 (a) illustrates the overall adapter tuning framework of our method. Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature transformation for task adaptation while maintaining the shape of feature maps consistent before and after each adaptation. Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen, leading to efficient task adaptation. The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategy [32, 2, 43].
Our D2ST-Adapter is designed independently from the backbone structure of the feature extractor, thus it can be readily integrated into most of classical pre-trained deep vision models. In particular, we instantiate the backbone of feature extractor with two dominant deep visual learning frameworks, namely ResNet [13] and ViT [7], respectively.
Figure 2: (a) Overall adapter tuning framework of our method. (b) DST-Adapter, a convolutional version of our model, which also follows dual-pathway architecture, whereas both pathways are constructed based on 3D convolution. (c) D2ST-Adapter, which is designed in a dual-pathway architecture and both pathways are modeled based on the proposed aDSTA illustrated in (d).
@@ 3.2 D2ST-Adapter
Dual-pathway Adapter Architecture for Spatio-Temporal Disentanglement. Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features. Most large vision models are typically pre-trained on image data, thus our D2ST-Adapter should be capable of capturing both the spatial features and temporal features. To this end, we design the D2ST-Adapter as a dual-pathway architecture shown in Figure 2 (c), in which the spatial pathway is responsible for capturing the spatial semantics while the temporal pathway focuses on learning the temporal dynamics. As a result, our model is able to encode the spatio-temporal features for video data in a disentangled manner.
@@
$\mathbf{F}_{i}^{\text{out}}=\text{GELU}\left(\mathcal{F}_{\text{S}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus\mathcal{F}_{\text{T}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\right)\cdot\mathbf{W}_{\text{up}}.$ (1)
Herein, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{C\times C^{\prime}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{C^{\prime}\times C}$ are the transformation matrices of two linear layers for downsampling and upsampling, respectively. $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ denote the disentangled feature adaptation by the spatial and temporal pathways respectively. Below we discuss two feasible modeling mechanisms for $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$, which leverages 3D convolution and our proposed Anisotropic Deformable Spatio-Temporal Attention, respectively.
@@
While it has been validated that leveraging 3D convolutions with different kernel shapes can learn the spatial and temporal features separately [8, 38], An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain. Constructing deep convolutional network by stacking convolutional layers can enlarge the receptive field whilst increasing the model size and incurring heavy computational burden, which is inconsistent with the lightweight design principle for adapters.
Modeling with Anisotropic Deformable Spatio-Temporal Attention. To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise anisotropic Deformable Spatio-Temporal Attention (aDSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency. It adapts the deformable attention [48] from 2D image space to 3D spatio-temporal space. As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling density in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions. As a result, aDSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
@@
$\mathbf{F}^{\prime}_{i}=\mathcal{F}_{\text{DPE}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}}).$ (2)
@@
$\mathcal{F}_{\text{tri-int}}(\mathbf{p})=\sum_{\mathbf{r}}g(p_{t},r_{t})\cdot g(p_{h},r_{h})\cdot g(p_{w},r_{w})\cdot\mathbf{F}^{\prime}_{i}(\mathbf{r}),$ (3)
@@
$\begin{split}&\mathbf{q}=\mathbf{F}^{\prime}_{i,(u_{t},u_{h},u_{w})}\cdot\mathbf{W}_{q},\mathbf{K}=\mathbf{P}\cdot\mathbf{W}_{k},\mathbf{V}=\mathbf{P}\cdot\mathbf{W}_{v},\\ &\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}=\text{softmax}(\mathbf{q}\mathbf{K}/\sqrt{C^{\prime}})\mathbf{V}\cdot\mathbf{W}_{o},\end{split}$ (4)
@@
Anisotropic Sampling Density. The learned offsets for reference points are restricted in a limited range for two reasons. First, it can prevent the potential training collapse that all reference points are shifted to the same point. Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region. As a result, more sampling of reference points typically leads to more powerful and fine-grained representations by the shifted points. The deformable attention samples reference points uniformly in the 2D feature space with the same sampling density along two spatial dimensions. In contrast, we configure anisotropic sampling density in the spatial and temporal domains. Consequently, our D2ST-Adapter can leverage aDSTA to model both spatial feature adaptation $\mathcal{F}_{\text{S}}$ in the spatial pathway and temporal feature adaptation $\mathcal{F}_{\text{T}}$ in the temporal pathway. Specifically, we define the ‘Sampling kernel’ for aDSTA as:
Definition 1
Sampling kernel: formulated as the sampling densities of reference points along the dimensions of time, height and width, respectively, denoted as $<n_{t},n_{s},n_{s}>$, s.t. $M=n_{t}\times n_{s}\times n_{s}$. $M$ is the total number of sampled reference points.
Intuitively, an aDSTA module whose sampling kernel has large $n_{t}$ and small $n_{s}$ are more capable of learning the temporal features than learning the spatial features. On the other hand, sampling more reference points in the spatial domain than the temporal domain ($n_{s}>n_{t}$) generally makes aDSTA focus on learning the spatial features. Thus, we can tailor aDSTA by configuring the sampling kernel to model $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ correspondingly. In particular, we tailor two versions of aDSTA:
  • aDSTA-S for modeling $\mathcal{F}_{\text{S}}$ in the spatial pathway, whose sampling kernel has relatively large $n_{s}$ and small $n_{t}$ ($n_{s}>n_{t}$).
@@
In practice, the values of $n_{t}$ and $n_{s}$ are tuned as hyper-parameters. Such modeling is analogous to classical property of 3D convolutional network that configuring the shape of the convolutional kernel can steer the convolutional operation to focus on learning features in different domain.

### v4

@@ Abstract
Adapting pre-trained image models to video modality has proven to be an effective strategy for robust few-shot action recognition. In this work, we explore the potential of adapter tuning in image-to-video model adaptation and propose a novel video adapter tuning framework, called Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter). It features a lightweight design, low adaptation overhead and powerful spatio-temporal feature adaptation capabilities. D2ST-Adapter is structured with an internal dual-pathway architecture that enables built-in disentangled encoding of spatial and temporal features within the adapter, seamlessly integrating into the single-stream feature learning framework of pre-trained image models. In particular, we develop an efficient yet effective implementation of the D2ST-Adapter, incorporating the specially devised anisotropic Deformable Spatio-Temporal Attention as its pivotal operation. This mechanism can be individually tailored for two pathways with anisotropic sampling densities along the spatial and temporal domains in 3D spatio-temporal space, enabling disentangled encoding of spatial and temporal features while maintaining a lightweight design. Extensive experiments by instantiating our method on both pre-trained ResNet and ViT demonstrate the superiority of our method over state-of-the-art methods. Our method is particularly well-suited to challenging scenarios where temporal dynamics are critical for action recognition. Code is available at https://github.com/qizhongtan/D2ST-Adapter.
footnotetext: * Equal contribution.   $\dagger$ Corresponding author.
Figure 1: Comparison with previous adapter-based methods. Both AIM and DUALPATH pre-learn the spatial and temporal features separately by duplicating the whole MSA modules, and insert vanilla adapter into each duplicated module for feature adaptation, which incurs heavy-weight model design. ST-Adapter designs the adapter for video data using 3D convolution whose performance is limited by the joint learning of spatio-temporal features. In contrast, our D2ST-Adapter is designed in a dual-pathway architecture to adapt the spatial and temporal features in a disentangled manner while maintaining the single-stream feature learning framework of the pre-trained model. Furthermore, we design the anisotropic Deformable Spatio-Temporal Attention (aDSTA), which configures anisotropic sampling densities in the spatial and temporal domains to model two pathways specifically and enables D2ST-Adapter to adapt features in a global view.
@@ 1 Introduction
Few-shot action recognition aims to learn an action recognition model from a set of base classes of video data, which can recognize novel categories of actions using only a few support samples. To this end, learning an effective feature extractor that is generalizable across different classes is crucial. A typical way [3, 45, 1, 33, 35] is to leverage pre-trained large image models [10, 29, 24] for feature learning by image-to-video model adaptation.
@@
Most existing methods [23, 34, 35] seek to adapt large vision models to few-shot action recognition by fine-tuning either the entire or partial model. While such methods have achieved promising performance, there are two important limitations that hamper them from exploiting the full potential of the pre-trained models. First, fully fine-tuning pre-trained models tends to overfit in few-shot learning scenarios, whereas partial fine-tuning is less effective due to its limited capacity for adaptation. Second, most large vision models are trained on image data due to the scarcity of video data. Thus, these fine-tuning based methods require to learn the temporal features as a post-processing step after feature extraction by the pre-trained backbone, typically by constructing an auxiliary module to model the temporal dynamics [28, 38, 33, 34], which has limited effectiveness compared to methods that incorporate temporal learning throughout the entire feature learning stage.
@@
As a prominent parameter-efficient fine-tuning technique [11, 50, 13, 9], adapter tuning [11, 4, 32] is particularly well-suited to few-shot learning scenarios due to much lower parameter-learning overhead. A straightforward way to apply adapter tuning to image-to-video model adaptation is to employ the vanilla image adapter [11] to adapt the pre-learned spatial and temporal features separately, which is exactly adopted by AIM [44] and DUALPATH [22]. As shown in Figure 1, both AIM and DUALPATH first pre-learn the spatial and temporal features separately by duplicating the multi-head self-attention modules, either in cascaded manner (AIM) or parallel manner (DUALPATH). Subsequently, they insert individual vanilla adapters in each duplicated module for specific feature adaptation. However, such methods inevitably result in a heavyweight model architecture and substantial computational overhead. As illustrated in Figure 1, ST-Adapter [21] is the first to design a specialized adapter for video data by incorporating a 3D convolution layer into the vanilla adapter to learn spatio-temporal features. Nevertheless, we investigate two potential limitations of ST-Adapter. First, it learns the spatial and temporal features jointly by 3D convolution, whereas it has been shown that in low-data scenarios, such a joint learning paradigm for spatio-temporal features in video data is inferior to learning the spatial and temporal features in a disentangled manner, such as SlowFast [6] and other two-stream designs [25, 31]. Second, the convolutional operation, especially with shallow layers for lightweight design, has a limited local receptive field in both the spatial and temporal domains, which also limits the performance of ST-Adapter.
In this work we present the Disentangled-and-Deformable Spatio-Temporal Adapter (D2ST-Adapter) to tackle the aforementioned limitations. Specifically, we design our D2ST-Adapter as a dual-pathway architecture, as illustrated in Figure 1, in which the spatial pathway is responsible for capturing the appearance features while the temporal pathway focuses on learning the temporal dynamics. This built-in disentangled encoding of spatio-temporal features equips our D2ST-Adapter with superior image-to-video model adaptation capabilities while maintaining a lightweight design and low adaptation overhead. As a result, it benefits from a more efficient model design compared to AIM and DUALPATH, which perform separate feature pre-learning and individual adaptation for spatial and temporal features by duplicating learning modules. Meanwhile, our D2ST-Adapter demonstrates superior feature learning capability for video data compared to ST-Adapter, owing to its disentangled adaptation scheme.
@@
We develop an efficient yet effective implementation of the D2ST-Adapter, featuring the specially devised anisotropic Deformable Spatio-Temporal Attention (aDSTA) as the pivotal operation to enable built-in disentangled encoding capability. It adapts the deformable attention [40] from 2D image space to 3D spatio-temporal space, which allows our D2ST-Adapter to encode features in both spatial and temporal domains in a global view while keeping lightweight modeling. One novel design of our aDSTA is that it can be tailored with anisotropic sampling densities along spatial and temporal domains, enabling specialized versions of aDSTA to model the spatial and temporal pathways separately. As shown in Figure 1, we tailor aDSTA-T with denser sampling along temporal domain than spatial domain for temporal pathway since it focuses on capturing temporal features. In contrast, the tailored aDSTA-S for spatial pathway samples denser points along spatial domain.
To conclude, we make the following contributions:
@@
  • We propose D2ST-Adapter, a novel video adapter tuning framework for few-shot action recognition. It is designed in a dual-pathway architecture featuring built-in disentangled adaptation for spatial and temporal features, enabling effective yet efficient image-to-video model adaptation for few shot action recognition.
@@
  • We devise aDSTA, which can be tailored with anisotropic sampling densities along spatial and temporal domains to model the spatial and temporal pathways separately, yielding a lightweight implementation of D2ST-Adapter.
@@ 2 Related Work
Few-shot Action Recognition. Most existing few-shot action recognition methods adopt the metric-based paradigm to classify videos, and primarily focus on two directions to deal with this task. The first one is to investigate the spatio-temporal modeling [52, 45, 28, 34, 43, 39, 36, 16]. STRM [28] enriches local patch features and global frame features for joint spatio-temporal modeling. MoLo [34] designs a motion autodecoder to explicitly extract motion dynamics in a unified network. Our method also aims to model spatio-temporal features effectively and efficiently based on adapter-tuning technique. Another direction focuses on designing effective metric learning strategies [3, 23, 33, 38, 12, 20]. OTAM [3] utilizes the DTW [19] algorithm to calculate video distances with strict temporal alignment. TRX [23] exhaustively enumerates all sub-sequences of support and query videos and matches them using attention mechanism. HyRSM [33] applies a novel bidirectional mean Hausdorff metric (referred to as Bi-MHM) to alleviate the strictly ordered constraints. We evaluate our D2ST-Adapter using the above three classical matching metrics to demonstrate its effectiveness and robustness.
Adapter Tuning. As a classical parameter-efficient fine-tuning method, adapter tuning is first proposed in [11], and quickly draws attention in many other research areas [4, 32, 9]. A common practice is to build up a lightweight module (named Adapter) which only consisting of negligible learnable parameters, and selectively plug it into a pre-trained model. During training, only the parameters of inserted adapters are updated while the original model remains frozen, leading to efficient task adaptation. Recently, some works apply this method to adapt image models for action recognition. AIM [44] duplicates multi-head self-attention modules and plugs adapters after them to separately learn the spatial and temporal features in a cascaded manner. Similarly, DUALPATH [22] explicitly builds two-stream architecture upon ViT and utilizes adapters to parallelly learn spatial and temporal features. ST-Adapter [21] employs depth-wise 3D convolution to construct adapter for feature adaptation, which endows it with spatio-temporal modeling capability. Our D2ST-Adapter follows the basic framework of ST-Adapter, and meanwhile optimize the essential technical designs for few-shot action recognition.
@@ 3.1 Overview
Problem Formulation of Few-shot Action Recognition. The task of few-shot action recognition typically follows episodic paradigm [30, 3, 33, 34]. An episode includes a support set $\mathcal{S}$ consisting of $N$ classes and $K$ labeled samples for each class (referred to as the $N$-way $K$-shot task), as well as a query set $\mathcal{Q}$ that contains unlabeled samples to be classified. In each episode, we aim to classify every query into one of the $N$ classes with the guidance of the support set. Such episodic task setting is consistently adopted during all the training, validation, and test stages.
Adapter Tuning Framework. We design a plug-and-play and lightweight adapter for video data, dubbed D2ST-Adapter, which can be integrated into most existing large image models. Thus we can leverage the powerful feature encoding capability of pre-trained models by efficient image-to-video model adaptation with only a small amount of parameter-tuning overhead. It is particularly well-suited to few-shot scenarios. Figure 2 (a) illustrates the overall adapter tuning framework of our method. Given a pre-trained feature extractor from a large model, our designed lightweight D2ST-Adapter can be selectively plugged into the middle layers of the feature extractor to perform feature adaptation. Only the inserted adapters are tuned during training while the pre-trained backbone keeps frozen. The learned features by such adapter tuning framework are further used for few-shot action recognition based on the metric-based strategies [3, 23, 33]. We instantiate the backbone of feature extractor with ResNet and ViT respectively.
Figure 2: (a) Overall adapter tuning framework of our method. (b) DST-Adapter, a convolutional version of our model, which follows dual-pathway architecture, whereas both pathways are constructed based on 3D convolution. (c) D2ST-Adapter, which is designed in a dual-pathway architecture and both pathways are modeled based on the proposed aDSTA illustrated in (d).
@@ 3.2 D2ST-Adapter
Dual-pathway Adapter Architecture. Successful action recognition from video data entails effective feature learning of both spatial semantic and temporal dynamic features. Most large vision models are typically pre-trained on image data, thus our D2ST-Adapter should be capable of capturing both the spatial features and temporal features. To this end, we design the D2ST-Adapter as a dual-pathway architecture shown in Figure 2 (c), in which the spatial pathway is responsible for capturing the spatial semantics while the temporal pathway focuses on learning the temporal dynamics. As a result, our model is able to encode the spatio-temporal features for video data in a disentangled manner.
@@
$\vspace{-3pt}\mathbf{F}_{i}^{\text{out}}=\text{GELU}\left(\mathcal{F}_{\text{S}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus\mathcal{F}_{\text{T}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\right)\cdot\mathbf{W}_{\text{up}}.$ (1)
Herein, $\mathbf{W}_{\text{down}}\in\mathbb{R}^{C\times C^{\prime}}$ and $\mathbf{W}_{\text{up}}\in\mathbb{R}^{C^{\prime}\times C}$ are the transformation matrices of two linear layers for downsampling and upsampling, respectively. $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ denote the disentangled feature adaptation by the spatial and temporal pathways respectively. Below we discuss two feasible modeling mechanisms for $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$, which leverages 3D convolution and our proposed aDSTA module, respectively.
@@
An important limitation of using 3D convolutions to construct D2ST-Adapter, which is also suffered by ST-Adapter, is that the convolutional operation has limited local receptive field either in the spatial or the temporal domain, especially with shallow layers for lightweight design.
Modeling with aDSTA. To address the limitation of the modeling mechanism based on 3D convolution (DST-Adapter), we devise anisotropic Deformable Spatio-Temporal Attention (aDSTA) which performs feature adaptation in a global view by sparse self-attention while maintaining high computational efficiency. It adapts the deformable attention [40] from 2D image space to 3D spatio-temporal space. As shown in Figure 2 (d), aDSTA first samples a group of reference points with anisotropic sampling densities in the spatio-temporal space and then learns the offset for each reference point to shift them to the more informative regions. As a result, aDSTA can learn a set of informative tokens which are further used as key and value pairs shared for all queries for sparse self-attention.
@@
$\vspace{-3pt}\mathbf{F}^{\prime}_{i}=\mathcal{F}_{\text{DPE}}(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}})\oplus(\mathbf{F}_{i}^{\text{in}}\cdot\mathbf{W}_{\text{down}}).$ (2)
@@
$\vspace{-5pt}\mathcal{F}_{\text{tri-int}}(\mathbf{p})=\sum_{\mathbf{r}}g(p_{t},r_{t})\cdot g(p_{h},r_{h})\cdot g(p_{w},r_{w})\cdot\mathbf{F}^{\prime}_{i}(\mathbf{r}),$ (3)
@@
$\vspace{-3pt}\begin{split}&\mathbf{q}=\mathbf{F}^{\prime}_{i,(u_{t},u_{h},u_{w})}\cdot\mathbf{W}_{q},\mathbf{K}=\mathbf{P}\cdot\mathbf{W}_{k},\mathbf{V}=\mathbf{P}\cdot\mathbf{W}_{v},\\ &\mathbf{Z}_{i,(u_{t},u_{h},u_{w})}=\text{softmax}(\mathbf{q}\mathbf{K}/\sqrt{C^{\prime}})\mathbf{V}\cdot\mathbf{W}_{o},\end{split}$ (4)
@@
Anisotropic sampling densities. The learned offsets for reference points are restricted in a limited range for two reasons. First, it can prevent the potential training collapse that all reference points are shifted to the same point. Second, such restriction can ensure that each reference point can be shifted to a unique location within a local region. As a result, denser sampling of reference points typically leads to finer-grained representations. The classical deformable attention samples reference points uniformly in 2D feature space with the same sampling densities along two spatial dimensions. In contrast, we configure anisotropic sampling densities in the spatial and temporal domains. Consequently, our D2ST-Adapter can leverage aDSTA to model both spatial feature adaptation $\mathcal{F}_{\text{S}}$ in the spatial pathway and temporal feature adaptation $\mathcal{F}_{\text{T}}$ in the temporal pathway.
We denote the sampling densities of reference points of aDSTA along the dimensions of time, height and width as $<n_{t},n_{s},n_{s}>$, thereby the total number of sampled reference points is $n_{t}\times n_{s}\times n_{s}$. Intuitively, an aDSTA module which samples large $n_{t}$ and small $n_{s}$ is more capable of learning the temporal features than learning the spatial features. On the other hand, sampling more reference points in the spatial domain than the temporal domain ($n_{s}>n_{t}$) generally makes aDSTA focus on learning the spatial features. Thus, we can tailor aDSTA by configuring the sampling densities to model $\mathcal{F}_{\text{S}}$ and $\mathcal{F}_{\text{T}}$ correspondingly:
  • aDSTA-S for modeling $\mathcal{F}_{\text{S}}$ in the spatial pathway, which is sampled with larger $n_{s}$ and smaller $n_{t}$ ($n_{s}>n_{t}$).
@@
The values of $n_{t}$ and $n_{s}$ are tuned as hyper-parameters.